# SOFIA/HAWC+ Far-Infrared Polarimetric Large Area CMZ Exploration Survey. V. The Magnetic Field Strength and Morphology in the Sagittarius C Complex

### v1

@@ Abstract
We present an analysis of the magnetic field strength and morphology in the Sagittarius C complex (Sgr C; G359.43-0.09) in the Milky Way Galaxy’s Central Molecular Zone (CMZ) using the 214 µm polarimetry data acquired with the High-Angular-Resolution Wideband Camera+ (HAWC+) instrument aboard the Stratospheric Observatory for Infrared Astronomy (SOFIA). We use several hundred magnetic field pseudovectors in the Sgr C region to trace the projected magnetic field orientation within cold molecular gas clouds, and as is the trend throughout the CMZ, they show a higher polarization fraction toward the periphery of the clouds. We conduct a modified Davis-Chandrasekhar-Fermi (DCF) analysis of individual clouds and find that the sky-plane magnetic field strength varies from highly turbulent regions having inferred strengths of $\sim 30~{}\mu{\rm G}$ to regions of relatively uniform field orientation having strengths of $\sim 1~{}{\rm mG}$. The magnetic field orientations suggest that outflows from active star-forming regions, such as the extended green object (EGO) G359.43-0.10 and the protostellar source FIR-4 (G359.43+0.02), cause high turbulence in their vicinity. The magnetic field direction is found to be tangential to the surface of the Sgr C H II region, as well as two [C II] emission cavities around this region. Several other features in the vicinity of Sgr C, especially numerous non-thermal filaments (NTFs) and a diffuse source of X-ray emission towards the southwest of the H II region, are discussed with regard to the observed magnetic field pseudovectors.
@@ 1 Introduction
The central half kiloparsec of the Milky Way disk, commonly referred to as the Central Molecular Zone (CMZ), is an ideal laboratory to study astrophysics in extreme environments (Morris & Serabyn, 1996; Ferrière et al., 2007; Bryant & Krabbe, 2021; Henshaw et al., 2023). Apart from the extensively studied star formation and Interstellar Medium (ISM) properties, a strong magnetic field has also been found to permeate the CMZ (Tsuboi et al., 1985; LaRosa et al., 2005; Morris, 2006; Ferrière, 2009; Morris, 2015; Akshaya & Hoang, 2024; Paré et al., 2024; Tress et al., 2024), especially in dense clouds (Yusef-Zadeh et al., 1999; Pillai et al., 2015; Hsieh et al., 2018; Butterfield et al., 2024a; Lu et al., 2024). The relatively strong magnetic field (0.1 – 1 mG) is said to counteract the gravitational collapse of clouds and contributes to the low star-formation rate in this region (Morris, 1993; Chuss et al., 2003; Kruijssen et al., 2014; Sormani et al., 2020; Morris, 2023).
@@
In this paper, we use the results of the Far-Infrared Polarimetric Large Area CMZ Exploration survey (FIREPLACE; PI: D. Chuss), a high-resolution dust polarimetry survey conducted at 214 µm, to study the magnetic field in the CMZ. As a common experimental probe of the sky-plane component of the magnetic field, dust polarimetry exploits the property that the spin axes of dust grains tend to be aligned with the magnetic field (Andersson et al., 2015). The measurement of polarized dust emission reveals the alignment of dust grains within these clouds, thereby indicating the sky-plane magnetic field component integrated along the line of sight (Chuss et al., 2003; Andersson et al., 2015).
@@
There are several possible mechanisms that could be responsible for the field orientations deduced from the FIREPLACE measurements. As the magnetic field lines are anchored inside the molecular clouds, the field lines could be sheared by the clouds’ internal motions (Parker, 1979). In a differentially orbiting medium like the CMZ, the shear stretches the magnetic field into a primarily toroidal geometry (Werner et al., 1988; Wardle & Konigl, 1990). The same effect can also be indicative of the direction of the internal shear within clouds (Aitken et al., 1991; Morris et al., 1992; Aitken et al., 1998; Novak et al., 2000; Chuss et al., 2003; Morris, 2015). Other candidates, such as cloud compression, which occurs at the interface of expanding H II regions or supernova blast waves, tend to shape the field to be parallel to the compressive interface.
@@
Indeed, a high-resolution magnetic field study, like the FIREPLACE survey, can benefit our understanding of many aspects of the CMZ, such as the magnetic field morphology within molecular clouds, and within portions of low-density ISM occupied by the NTFs (Yusef-Zadeh et al., 1984; Morris & Yusef-Zadeh, 1985; Heywood et al., 2022; Yusef-Zadeh et al., 2022a). These NTFs are manifested as spatially extended synchrotron emission from relativistic electrons constrained to diffuse along a magnetic flux tube. The NTFs within the CMZ are generally oriented perpendicular to the Galactic plane, which implies a large-scale vertical magnetic field pervading the Galactic Centre (GC), thereby pointing to the existence of an organized inter-cloud magnetic field (Morris & Serabyn, 1996; Heywood et al., 2022). Furthermore, Serabyn & Morris (1994) argued that turbulent instabilities at the ${\rm H_{2}}$/H II interface could mix two orthogonal magnetic field systems inside and outside that interface, which facilitates magnetic field line reconnection, which in turn can boost free electrons in the ISM to relativistic velocities, thereby forming NTFs as those electrons diffuse along the magnetic field lines in the inter-cloud ISM (Serabyn & Morris, 1994; Morris, 1996; Chuss et al., 2003). In addition, many other hypotheses have been offered to account for the NTFs (references in Morris, 1996; Boldyrev & Yusef-Zadeh, 2006; Yusef-Zadeh & Königl, 2004; Banda-Barragán et al., 2016, 2018; Yusef-Zadeh & Wardle, 2019). For example, an alternative hypothesis for generating relativistic electrons that does not directly involve magnetic fields invokes diffusive shock acceleration where stellar winds from massive young stars impact relatively dense interstellar material (Rosner & Bodo, 1996).
@@
Previous studies of the GC magnetic field, including those using the Submillimeter Polarimeter for Antarctic Remote Observations (SPARO; Novak et al. 2003), Planck (Planck Collaboration et al., 2015), the Polarized Instrument for the Long-wavelength Observation of the Tenuous ISM (PILOT) (Mangilli et al., 2019), and the Atacama Cosmology Telescope (ACT; Guan et al. 2021), have focused on the large-scale magnetic field, with typical resolutions $\gtrsim 1\arcmin{}$. These works reported a largely uniform magnetic field that permeates the CMZ, and is oriented at a $\sim 20$ angle with respect to the Galactic plane. However, conclusions made from these surveys remain on a global level, since the spatial resolution of these surveys does not generally resolve the internal structure of individual molecular clouds. High-resolution surveys, such as those reported by Chuss et al. (2003) and Lu et al. (2024) found a correlation between the magnetic field orientation and molecular cloud density, though they do not have enough coverage to cover a sufficiently large portion of the CMZ to draw global conclusions. Along the same line, Bally et al. (2024) studied the magnetic field inside the prominent H II region using the JWST-NIRCam images. With the FIREPLACE survey dataset, a recent study by Paré et al. (2025) (or FIREPLACE IV) demonstrated the advantage of studying the magnetic field orientation in the CMZ using datasets with both high resolution and large coverage. In this paper, we focus on studying the magnetic field within one of its sub-regions, the Sgr C complex.
@@
Sgr C was initially reported as one of the most prominent H II regions in the GC (Downes & Maxwell, 1966), and it was later detailed by Liszt (1985) and Liszt & Spiker (1995), who noted the presence of a prominent NTF apparently connected to the H II region. Figure 1 portrays this region at multiple wavelengths overlaid with a line integral contour image (LIC; Cabral & Leedom 1993) to represent the magnetic field direction made by measurements reported by Paré et al. (2024) (or FIREPLACE III).
@@
To highlight important structures, Figure 2 shows a simplified schematic diagram of Sgr C.
Figure 2: Schematic diagram of the main features within the Sgr C complex on the sky-plane, including FIR-4 (blue ellipse), the major molecular clouds detected in far-IR (purple clouds), the Sgr C H II region and NTF detected in radio continuum (yellow circle and long, vertical rectangle, respectively), the G359.44-0.10 EGO (Kendrew et al. 2013; green star), and the radio and mid-IR structure source C (yellow cloud). The velocities of the molecular clouds depicted in magenta are obtained using the molecular line observation of the 3 mm Mopra survey reported by Jones et al. (2012), detailed in Section 4.2.
@@
In this paper, we examine the magnetic field structure of the Sgr C complex using the FIREPLACE survey. The paper is organized as follows. In Section 2, we briefly describe the survey data used. In Section 3, we present analyses of the Sgr C magnetic field. In Section 4, we use multi-wavelength observations to investigate numerous intriguing structures in the Sgr C region where the magnetic field has apparently become manifest. In Section 5, we present a detailed examination of several features in which the magnetic field is likely playing an important role. We finally summarize our conclusions in Section 6. Throughout this paper, we adopt 8.2 kpc as the distance to the GC (GRAVITY Collaboration et al., 2019).
@@ 3.1 Overview
Figure 3: The FIREPLACE magnetic field pseudovectors in the Sgr C complex, displayed with the 214 µm intensity image in the background, rendered using measurements of Paré et al. (2024). The yellow dashes are the 214 µm magnetic field pseudovectors made using Nyquist sampling at the HAWC+ beamsize, that fulfill the selection criteria described in Section 2. The length of the vectors is scaled by their polarization fraction. The legend in the lower right corner denotes the length of a 5% polarized pseudovector. The magenta contours are made from the MeerKAT/L-band observation of Heywood et al. (2022) at levels 0.0004 and 0.002 Jy/beam, with the brighter contour indicating higher intensity level. The orange dashes are the JCMT/POL2 magnetic field detections at 850 µm (Lu et al., 2024). For visual clarity, we display all 850 µm magnetic field detections with polarization fraction exceeding $p=0.08$ with a length equivalent to an 8% polarized pseudovector.
A similar pseudovector representation of the magnetic field measurements is adopted throughout this paper. The 214 µm emission is dominated by thermal emission from dust in molecular clouds, including the $-90$ km s-1 cloud, the $-65$ km s-1 cloud in the south, and FIR-4. Throughout the extended region surrounding Sgr C, the 214 µm polarization measurements that satisfy the selection criteria are clearly present with high significance throughout the far-IR-bright emission region. Measurements in faint regions do not commonly pass the selection criteria, owing either to low signal-to-noise ratio in either or both total intensity or polarization fractions, as outlined in Section 2. Pseudovectors in faint regions can be contaminated by residual systematic effects and/or confused by unrelated emissions originating outside the CMZ (Chuss et al. 2019; Butterfield et al. 2024a). However, in our field of view, our conservative selection criteria have largely mitigated such issues.
Throughout this paper, we compare our 214 µm magnetic field orientations to the 850 µm JCMT magnetic field orientations reported by Lu et al. (2024), depicted as orange in Figure 3. The magnetic field measurements of Lu et al. (2024) were obtained with a spatial resolution similar to that in Paré et al. (2024) and were produced under a similar set of cutoff thresholds. From Figure 3, the 850 µm magnetic field orientations do not generally agree with those at 214 µm. It is likely that, given the large difference in the survey wavelengths, the two studies sampled different temperature regimes along the line of sight.
@@
In particular, we observe an elongated far-IR-dark region along $l=0.08$ in Figure 3. This region corresponds to the Galactic plane and is home to many compact mid-IR sources and short and curved radio filamentary structures (see Panel #2 of Figure 7; Liszt & Spiker 1995; Carey et al. 2009; Yusef-Zadeh et al. 2009; Hankins et al. 2020; Cotera et al. 2024). The low 214 µm intensity and polarization non-detections there indicate a relative absence of cold dust, where it has either been heated by the local radiation field, expelled by stellar winds and expanding H II regions, or there is simply a dearth of material there.
@@ 3.2 DCF Analysis
where $\rho_{m}$ is the mass density, $\sigma_{v}$ is the velocity dispersion of the gas content, and $\sigma_{\phi}$ is the angular dispersion of the magnetic field pseudovectors. The DCF method is based on the assumption that the energy density of the turbulent or random gas motions is equal to the energy density of the fluctuations of the magnetic field about the local mean field, where its velocity dispersion is assumed to be isotropic. Though the general principle of the DCF method is commonly agreed upon, specific measurements of the input parameters of Equation 2 are under active refinement and debate, especially in the context of a dynamically complex region like the GC (Guerra et al. 2023; Butterfield et al. 2024b; Lu et al. 2024). We address our methodology for estimating the inputs $\rho_{m}$, $\sigma_{v}$, and $\sigma_{\phi}$ in the following subsections. In our DCF analysis, we do not consider the effect of shear on the sky-plane magnetic field strength $|\mathrm{\bf B}_{\rm POS}|$ (e.g. the shear-flow approximation considered by Guerra et al. 2023). Therefore, there is a caveat that our analysis can overestimate $|\mathrm{\bf B}_{\rm POS}|$ as the shear could lower the angular dispersion. However, due to the highly turbulent nature of the Sgr C region, shear does not play a significant role in the estimated magnetic field strength, as it is only significant in clouds with uniform motion.

### v2

@@ Abstract
We present an analysis of the magnetic field strength and morphology in the Sagittarius C complex (Sgr C; G359.43-0.09) in the Milky Way Galaxy’s Central Molecular Zone (CMZ), using the 214 µm polarimetry data acquired with the High-resolution Airborne Wide-band Camera (HAWC+) instrument aboard the Stratospheric Observatory for Infrared Astronomy (SOFIA). We conduct a modified Davis-Chandrasekhar-Fermi (DCF) analysis of individual clouds and find that the sky-plane magnetic field strength varies from highly turbulent regions having inferred strengths of $\sim 30~{}\mu{\rm G}$ to regions of relatively uniform field orientation having strengths of $\sim 300~{}\mu{\rm G}$. Several hundred magnetic field pseudovectors in the Sgr C region were measured to trace the projected magnetic field orientation within cold molecular clouds, and as is the trend throughout the CMZ, they show a higher polarization fraction toward the periphery of the clouds. The magnetic field orientations suggest that outflows from active star-forming regions, such as the G359.43-0.10 extended green object (EGO) and the protostellar source FIR-4 (G359.43+0.02), cause high turbulence in their vicinity. The magnetic field direction is found to be tangential to the surface of the Sgr C H II region, which displays spatial correspondence with two [C II] emission cavities reported in the H II region, signifying a compression front between the H II region and the surrounding dense clouds. Several other features in the vicinity of Sgr C, especially numerous non-thermal radio filaments (NTFs) and a diffuse source of X-ray emission to the immediate southwest of the H II region, are discussed with regard to the magnetic field measurements.
@@ 1 Introduction
The central half kiloparsec of the Milky Way disk, commonly referred to as the Central Molecular Zone (CMZ), is an ideal laboratory to study astrophysics in extreme environments given its high density, massive star-formation, and complex structure (see Morris & Serabyn, 1996; Ferrière et al., 2007; Bryant & Krabbe, 2021; Henshaw et al., 2023, and references within). Apart from the extensively studied star formation and Interstellar Medium (ISM) properties, a strong magnetic field has also been found to permeate the CMZ (Tsuboi et al., 1985; LaRosa et al., 2005; Morris, 2006; Ferrière, 2009; Morris, 2015; Akshaya & Hoang, 2024; Paré et al., 2024; Tress et al., 2024), especially in dense clouds (Yusef-Zadeh et al., 1999; Pillai et al., 2015; Hsieh et al., 2018; Butterfield et al., 2024a; Lu et al., 2024). The relatively strong magnetic field (0.1 – 1 mG) is said to counteract the gravitational collapse of clouds and contributes to the lower than expected star-formation rate in this region given its high density (Morris, 1993; Chuss et al., 2003; Kruijssen et al., 2014; Morris, 2023), therefore of great interest to be studied.
@@
In this paper, we use the results of the Far-Infrared Polarimetric Large Area CMZ Exploration survey (FIREPLACE; PI: D. Chuss), a high-resolution dust polarimetry survey conducted at 214 µm, to study the magnetic field in the CMZ. As a common experimental probe of the sky-plane component of the magnetic field, dust polarimetry exploits the property that the spin axes of dust grains tend to be aligned with the magnetic field (Andersson et al., 2015). The measurement of polarized emission from clouds therefore reveals the alignment of dust grains within, indicating the sky-plane magnetic field component integrated along the line of sight (Chuss et al., 2003; Andersson et al., 2015).
@@
There are several possible mechanisms that could be responsible for the field orientations deduced from dust polarimetry observations. As the magnetic field lines are anchored inside the molecular clouds, the field lines could be sheared by the clouds’ internal motions (Parker, 1979). In a differentially orbiting medium like the CMZ, the shear stretches the magnetic field into a primarily toroidal geometry observed in large scale (Werner et al., 1988; Wardle & Konigl, 1990). The same effect can also be indicative of the direction of the internal shear within clouds on a smaller scale (Aitken et al., 1991; Morris et al., 1992; Aitken et al., 1998; Novak et al., 2000; Chuss et al., 2003; Morris, 2015). Other candidates that can contribute to the magnetic field orientations, such as cloud compression, which occurs at the interface of expanding H II regions or supernova blast waves, tend to shape the field to be parallel to the compressive interface.
@@
Indeed, a high-resolution magnetic field study, like the FIREPLACE survey, can benefit our understanding of many aspects of the CMZ, such as the magnetic field morphology within molecular clouds and within portions of low-density ISM occupied by the NTFs (Yusef-Zadeh et al., 1984; Morris & Yusef-Zadeh, 1985; Heywood et al., 2022; Yusef-Zadeh et al., 2022a). These NTFs are manifested as spatially extended synchrotron emission from relativistic electrons constrained to diffuse along a magnetic flux tube. The NTFs within the CMZ are generally oriented perpendicular to the Galactic plane, which implies a large-scale vertical magnetic field pervading the Galactic Centre (GC), thereby pointing to the existence of an organized inter-cloud magnetic field (Morris & Serabyn, 1996; Heywood et al., 2022). Furthermore, Serabyn & Morris (1994) argued that turbulent instabilities at the ${\rm H_{2}}$/H II interface could mix two orthogonal magnetic field systems inside and outside that interface, which facilitates magnetic field line reconnection. Such phenomenon can boost free electrons in the ISM to relativistic velocities, thereby forming NTFs as those electrons diffuse along the magnetic field lines in the inter-cloud ISM (Serabyn & Morris, 1994; Morris, 1996; Chuss et al., 2003). In addition, many other hypotheses have been offered to account for the NTFs (see references in Morris, 1996; Boldyrev & Yusef-Zadeh, 2006; Yusef-Zadeh & Königl, 2004; Banda-Barragán et al., 2016, 2018; Yusef-Zadeh & Wardle, 2019). For example, an alternative hypothesis for generating relativistic electrons that does not directly involve magnetic fields invokes diffusive shock acceleration where stellar winds from massive young stars impact relatively dense interstellar material (Rosner & Bodo, 1996).
@@
Previous studies of the GC magnetic field, including those using the Submillimeter Polarimeter for Antarctic Remote Observations (SPARO; Novak et al. 2003), Planck (Planck Collaboration et al., 2015), the Polarized Instrument for the Long-wavelength Observation of the Tenuous ISM (PILOT) (Mangilli et al., 2019), and the Atacama Cosmology Telescope (ACT; Guan et al. 2021) have focused on the large-scale magnetic field, with typical resolutions $\gtrsim 1\arcmin{}$. These works reported a largely uniform magnetic field that permeates the CMZ, and is oriented at a $\sim 20$ angle with respect to the Galactic plane. However, conclusions made from these surveys remain on a global level, since the spatial resolution of these surveys does not generally resolve the internal structure of individual molecular clouds. High-resolution surveys conducted on subregions within the CMZ, such as those reported by Chuss et al. (2003) and Lu et al. (2024) found a correlation between the magnetic field orientation and molecular cloud density, though they do not have enough coverage to cover a sufficiently large portion of the CMZ to draw global conclusions. Along the same line in Sgr C, Bally et al. (2024) studied the magnetic field inside the prominent H II region using the JWST-NIRCam images. With the FIREPLACE survey dataset, a recent study by Paré et al. (2025) (or FIREPLACE IV) demonstrated the advantage of studying the magnetic field orientation in the CMZ using datasets with both high resolution and large coverage.
@@
In this paper, we focus on studying the magnetic field within one of the sub-regions in the CMZ — the Sgr C complex. Sgr C was initially reported as one of the most prominent H II regions in the GC (Downes & Maxwell, 1966), and it was later detailed by Liszt (1985) and Liszt & Spiker (1995), who noted the presence of a prominent NTF apparently connected to the H II region. Figure 1 portrays this region at multiple wavelengths overlaid with a line integral contour (LIC; Cabral & Leedom 1993) image to represent the magnetic field direction made by measurements reported by Paré et al. (2024) (or FIREPLACE III).
@@
Figure 2 shows a simplified schematic diagram of Sgr C to highlight the important structures within.
Figure 2: Schematic diagram of the main features within the Sgr C complex on the sky-plane, including FIR-4 (blue ellipse), the major molecular clouds detected in far-IR (purple clouds denoted by their measured velocities), the Sgr C H II region and NTF detected in radio continuum (yellow circle and long, vertical rectangle, respectively), the G359.44-0.10 EGO (Kendrew et al. 2013; green star), and the radio and mid-IR structure known as source C (yellow cloud). The velocities of the molecular clouds depicted in magenta are obtained using the molecular line observation of the 3 mm Mopra survey reported by Jones et al. (2012), detailed in Section 4.2.
@@
In this paper, we study the magnetic field measurements from the FIREPLACE survey within the Sgr C complex, in the context of other known structures in this region. The paper is organized as follows. In Section 2, we briefly describe the FIREPLACE survey. In Section 3, we present analyses of the Sgr C magnetic field measured in 214 µm. In Section 4, we use multi-wavelength observations to investigate numerous intriguing structures in the Sgr C region where the magnetic field has apparently become manifest. In Section 5, we present a detailed examination of several features in which the magnetic field is likely playing an important role. We finally summarize our conclusions in Section 6. Throughout this paper, we adopt 8.2 kpc as the distance to the GC (GRAVITY Collaboration et al., 2019).
@@ 3.1 Overview
Figure 3: The FIREPLACE magnetic field pseudovectors in the Sgr C complex, displayed with the 214 µm intensity image in the background, rendered using measurements of Paré et al. (2024). The yellow dashes are the 214 µm magnetic field pseudovectors made using Nyquist sampling at the HAWC+ beamsize, that fulfill the selection criteria described in Section 2. The length of the vectors is scaled by their polarization fraction. The legend in the lower right corner denotes the length of a 5% polarized pseudovector. The magenta contours are made from the MeerKAT/L-band observation of Heywood et al. (2022) at levels 0.0004 and 0.002 Jy/beam, with the brighter contour indicating higher intensity level. The orange dashes are the JCMT/POL2 magnetic field detections at 850 µm (Lu et al., 2024). The JCMT/POL2 magnetic field measurement rendition share the same scaling relation with the FIREPLACE scale according to their polarization fraction. For visual clarity, we display all 850 µm magnetic field detections with polarization fraction exceeding $p=0.08$ with a length equivalent to an 8% polarized pseudovector.
A similar pseudovector representation of the magnetic field measurements is adopted throughout this paper. The 214 µm emission is dominated by thermal emission from dust in molecular clouds, including the $-90$ km s-1 cloud in the north, the $-65$ km s-1 cloud in the south, and FIR-4. Throughout the extended region surrounding Sgr C, the 214 µm polarization measurements that satisfy the selection criteria are clearly present with high significance throughout the far-IR-bright emission region. Measurements in faint regions do not commonly pass the selection criteria, owing either to low signal-to-noise ratio in either or both total intensity or polarization fractions, as outlined in Section 2. Pseudovectors in faint regions can be contaminated by residual systematic effects and/or confused by unrelated emissions originating outside the CMZ (Chuss et al. 2019; Butterfield et al. 2024a). However, in our field of view, our conservative selection criteria have largely mitigated such issues.
Throughout this paper, we compare our 214 µm magnetic field orientations to the 850 µm JCMT magnetic field orientations reported by Lu et al. (2024), depicted as orange in Figure 3. The magnetic field measurements of Lu et al. (2024) were obtained with a spatial resolution similar to that in Paré et al. (2024) and were produced under a similar set of cutoff thresholds. From Figure 3, the 850 µm magnetic field orientations do not generally agree with those at 214 µm. It is likely that, given the large difference in the survey wavelengths, these two studies sample different regimes of dust temperature along the line of sight (see, e.g., the dust temperature study reported by Hankins et al., 2017).
@@
In particular, we observe an elongated far-IR-dark region along $b=-0.08$ in Figure 3. This region corresponds to the Galactic plane and is home to many compact mid-IR sources and short and curved radio filamentary structures (see Panel #2 of Figure 7; Liszt & Spiker 1995; Carey et al. 2009; Yusef-Zadeh et al. 2009; Hankins et al. 2020; Cotera et al. 2024). The low 214 µm intensity and polarization non-detections there indicate a relative absence of cold dust, where it has either been heated by the local radiation field, expelled by stellar winds and expanding H II regions, or there is simply a dearth of material there.
@@ 3.2 DCF Analysis
where $\rho_{m}$ is the mass density, $\sigma_{v}$ is the velocity dispersion of the gas content, and $\sigma_{\phi}$ is the angular dispersion of the magnetic field pseudovectors. The DCF method is based on the assumption that the energy density of the turbulent or random gas motions is equal to the energy density of the fluctuations of the magnetic field about the local mean field, where its velocity dispersion is assumed to be isotropic. Though the general principle of the DCF method is commonly agreed upon, specific measurements of the input parameters of Equation 2 are under active refinement and debate, especially in the context of a dynamically complex region like the GC (Guerra et al. 2023; Butterfield et al. 2024b; Lu et al. 2024). In our DCF analysis, we do not consider the effect of shear on the sky-plane magnetic field strength $|\mathrm{\bf B}_{\rm POS}|$ (e.g. the shear-flow approximation considered by Guerra et al. 2023). Therefore, there is a caveat that our analysis can overestimate $|\mathrm{\bf B}_{\rm POS}|$ as the shear could lower the angular dispersion. However, due to the highly turbulent nature of the Sgr C region, shear does not play a significant role in the estimated magnetic field strength, as it is only significant in clouds with uniform motion. We address our methodology for estimating the inputs $\rho_{m}$, $\sigma_{v}$, and $\sigma_{\phi}$ to calculated the magnetic field strength $|\mathrm{\bf B}_{\rm POS}|$ as in Equation 2 in the following subsections.